# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 6. Lower Bound for Large Derivations
Tightness of $\{\bar{\theta}_{\varepsilon,m}\}$ in $\mathbb{X}_{\alpha,\beta}$ follows by the same method as in the proof of Proposition 3.7, thus we omit the proof. According to Jakubowski-Skorokhod representation theorem [Jak97], for any sequence $\varepsilon_{k}\rightarrow 0$, there exists a subsequence $(\varepsilon_{k},m_{k})$ (still denoted by $(\varepsilon,m)$), a new stochastic basis (still denoted by $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\in[0,T]},\mathbb{P})$), a cylindrical Wiener process (still denoted by $W$), a sequence of $\mathbb{X}_{\alpha,\beta}$-valued random elements with the same laws as $\{\bar{\theta}_{\varepsilon,m}\}$ on the new probability space (still denoted by $\{\bar{\theta}_{\varepsilon,m}\}$) and a limit point $\bar{\theta}$, such that $\mathbb{P}$-almost surely,
$\bar{\theta}_{\varepsilon,m}\rightarrow\bar{\theta}\,\text{ in }\mathbb{X}_{\alpha,\beta}.$
We claim that $\bar{\theta}$ is a weak solution of the (1.1). For any $\varphi\in C^{\infty}\left([0,T];H^{\infty}(\mathbb{T}^{2})\right)$ and $t\in[0,T]$, we will pass to the limits in all terms of
@@
$\left\langle\bar{\theta}_{\varepsilon,m}(t),\varphi(t)\right\rangle$ $+\int_{0}^{t}\left\langle\bar{\theta}_{\varepsilon,m},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s=\left\langle\bar{\theta}_{\varepsilon,m}(0),\varphi(0)\right\rangle+\int_{0}^{t}\langle\bar{\theta}_{\varepsilon,m},\partial_{s}\varphi\rangle\,\mathrm{d}s$ (6.6)
@@
$-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,\mathrm{d}s$
@@
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi,P_{m}\,\mathrm{d}W(s)\rangle+\int_{0}^{t}\left\langle P_{m}g,\Lambda^{2\beta}\varphi\right\rangle\,\mathrm{d}s.$
An analysis similar to that in the proof of Lemma 4.4 shows that, $\mathbb{P}$-almost surely, for all $t\in[0,T],$ as $\varepsilon\rightarrow 0,$
$\left\langle\bar{\theta}_{\varepsilon,m}(t),\varphi(t)\right\rangle$ $\rightarrow\left\langle\bar{\theta}(t),\varphi(t)\right\rangle,\quad\int_{0}^{t}\left\langle\bar{\theta}_{\varepsilon,m},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s\rightarrow\int_{0}^{t}\left\langle\bar{\theta},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s,$
$\left\langle\bar{\theta}_{\varepsilon,m}(0),\varphi(0)\right\rangle$ $\rightarrow\left\langle\bar{\theta}(0),\varphi(0)\right\rangle,\quad\int_{0}^{t}\langle\bar{\theta}_{\varepsilon,m},\partial_{s}\varphi\rangle\,\mathrm{d}s\rightarrow\int_{0}^{t}\langle\bar{\theta},\partial_{s}\varphi\rangle\mathrm{d}s,$
and
$\int^{t}_{0}\left\langle R_{2}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,$ $\mathrm{d}s-\int_{0}^{t}\left\langle R_{1}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,\mathrm{d}s\rightarrow$
$-\int^{t}_{0}\left\langle R_{2}\bar{\theta},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle R_{1}\bar{\theta},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}\right\rangle\,\mathrm{d}s.$
It remains to show the convergence of the last two terms in (6.6). According to Doob’s $L^{2}$ inequality and Itô isometry, it holds that
@@
$\mathbb{E}\left(\sup_{0\leqslant t\leqslant T}\left|\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),P_{m}\,\mathrm{d}W(s)\rangle\right|^{2}\right)\lesssim\varepsilon\left\|\Lambda^{2\beta}\varphi\right\|_{L^{2}([0,T];H)}^{2}\rightarrow 0.$
Thus we can pass to a further subsequence such that, $\mathbb{P}$-almost surely,
@@
$\sup_{0\leqslant t\leqslant T}\left|\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),P_{m}\,\mathrm{d}W(s)\rangle\right|\rightarrow 0.$
Finally, Hölder’s inequality implies that
$\left|\int_{0}^{t}\left\langle P_{m}g-g,\Lambda^{2\beta}\varphi\right\rangle\,\mathrm{d}s\right|\lesssim\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}\left\|P_{m}g-g\right\|_{L^{2}([0,T];H)}\rightarrow 0.$
Therefore we conclude that, $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
@@
Step 2. $\bar{\theta}$ satisfies $H^{\alpha-2\beta}$-energy inequality.
Applying Itô’s formula to $\bar{\theta}_{\varepsilon,m}$, we deduce that $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\frac{1}{2}$ $\left\|\bar{\theta}_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ (6.7)
$+\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle+\frac{\varepsilon}{2}\left\|P_{m}\Lambda^{\alpha}\right\|_{HS}^{2}t+\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}(s),P_{m}g(s)\right\rangle\,\mathrm{d}s.$
Letting $\varepsilon\rightarrow 0$ in (6.7), it follows from the fact that $\bar{\theta}_{\varepsilon,m}\rightarrow\bar{\theta}\,\text{ in }\mathbb{X}_{\alpha,\beta}$ and the lower semi-continuity of the $H^{s}(\mathbb{T}^{2})$ norms that
$\frac{1}{2}\|\bar{\theta}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\bar{\theta}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\liminf_{\varepsilon\rightarrow 0}\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
Let $M^{\varepsilon,m}_{3}(t):=\int_{0}^{t}\sqrt{\varepsilon}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}(s),P_{m}\,\mathrm{d}W(s)\right\rangle$. Applying Doob’s $L^{2}$ inequality, we obtain
$\mathbb{E}\left(\sup_{t\leqslant T}\left|M_{3}^{\varepsilon,m}(t)\right|^{2}\right)\leqslant 4\mathbb{E}\left\langle\langle M_{3}^{\varepsilon}\right\rangle\rangle_{T}\leqslant 4\varepsilon\mathbb{E}\int_{0}^{T}\left\|\bar{\theta}_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\rightarrow 0.$
Hence we can pass to a further subsequence so that $M_{3}^{\varepsilon,m}(t)\rightarrow 0$ almost surely. By the scaling regime (2.8), the correction term $\frac{\varepsilon}{2}\left\|P_{m}\Lambda^{\alpha}\right\|_{HS}^{2}t\rightarrow 0$. Finally, according to the weak convergence $\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}\rightarrow\Lambda^{2\alpha-2\beta}\bar{\theta}$ in $L^{2}([0,T];H)$ and the strong convergence $P_{m}g\rightarrow g$ in $L^{2}([0,T];H)$, it holds that
$\left|\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m},P_{m}g\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta},g\right\rangle\,\mathrm{d}s\right|\rightarrow 0.$
@@
Step 3. $\bar{\theta}=\theta$ in $\mathbb{X}_{\alpha,\beta}.$
Thanks to the choice $\theta\in\mathcal{C}_{0}$, it follows from Definition 6.1 that $\theta=\bar{\theta}$ in $\mathbb{X}_{\alpha,\beta}$. Hence $\pi^{\varepsilon,m,\theta}$ converges to $\delta_{\theta}$ in law on $\mathbb{X}_{\alpha,\beta}$ as $\varepsilon\rightarrow 0$. ∎
@@
For every $\varepsilon,\delta>0$ and $m\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m}$ be a solution of (3.7) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$. Suppose that $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1).
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the law of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
@@
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.8) holds for $(\varepsilon,m)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$, the restricted lower bound (6.8) holds as well.
@@
Proof of Proposition 6.3.
For any open set $G\subset\mathbb{X}_{\alpha,\beta}$ and any $\eta>0$, there exists an element $\theta^{0}\in G\cap\mathcal{C}$ such that
@@
Using (6.9), (6.10) and the arbitrariness of $\eta$, we obtain that

### v3

@@ 6. Lower Bound for Large Deviations
Step 1. $\bar{\theta}$ is a weak solution of the skeleton equation. Tightness of $\{\bar{\theta}_{\varepsilon,m(\varepsilon)}\}$ in $\mathbb{X}_{\alpha,\beta}$ follows by the same method as in the proof of Proposition 3.6, thus we omit the proof. According to Jakubowski-Skorokhod representation theorem [Jak97], for any sequence $\varepsilon_{k}\rightarrow 0$, there exists a subsequence $(\varepsilon_{k},m_{k}(\varepsilon))$ (still denoted by $(\varepsilon,m(\varepsilon))$), a new stochastic basis (still denoted by $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\in[0,T]},\mathbb{P})$), a cylindrical Wiener process (still denoted by $W$), a sequence of $\mathbb{X}_{\alpha,\beta}$-valued random elements with the same laws as $\{\bar{\theta}_{\varepsilon,m(\varepsilon)}\}$ on the new probability space (still denoted by $\{\bar{\theta}_{\varepsilon,m(\varepsilon)}\}$), and a limit point $\bar{\theta}$, such that $\mathbb{P}$-almost surely, $\bar{\theta}_{\varepsilon,m(\varepsilon)}\rightarrow\bar{\theta}\,\text{ in }\mathbb{X}_{\alpha,\beta}$. We claim that $\bar{\theta}$ is a weak solution of (1.1). For any $\varphi\in C^{\infty}\left([0,T];H^{\infty}(\mathbb{T}^{2})\right)$ and $t\in[0,T]$, we will pass to the limits in all terms of
@@
$\left\langle\bar{\theta}_{\varepsilon,m(\varepsilon)}(t),\varphi(t)\right\rangle+\int_{0}^{t}\left\langle\bar{\theta}_{\varepsilon,m(\varepsilon)},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s=\left\langle\bar{\theta}_{\varepsilon,m(\varepsilon)}(0),\varphi(0)\right\rangle+\int_{0}^{t}\langle\bar{\theta}_{\varepsilon,m(\varepsilon)},\partial_{s}\varphi\rangle\,\mathrm{d}s$ (6.5)
@@
$-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\bar{\theta}_{\varepsilon,m(\varepsilon)},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m(\varepsilon)}\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\bar{\theta}_{\varepsilon,m(\varepsilon)},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m(\varepsilon)}\right\rangle\,\mathrm{d}s$
@@
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi,P_{m(\varepsilon)}\,\mathrm{d}W(s)\rangle+\int_{0}^{t}\left\langle P_{m(\varepsilon)}g,\Lambda^{2\beta}\varphi\right\rangle\,\mathrm{d}s.$
It follows from the proof of Lemma 4.4 that, $\mathbb{P}$-almost surely, for all $t\in[0,T],$ as $\varepsilon\rightarrow 0,$ the first six terms in (6.5) converge. For the stochastic integral term, applying Doob’s $L^{2}$-inequality, we obtain
@@
$\mathbb{E}\left(\sup_{t\in[0,T]}\left|\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi,P_{m(\varepsilon)}\,\mathrm{d}W(s)\rangle\right|^{2}\right)\leqslant 4\varepsilon\mathbb{E}\left|\int_{0}^{T}\langle\Lambda^{2\beta}\varphi,P_{m(\varepsilon)}\,\mathrm{d}W(s)\rangle\right|^{2}.$
According to Itô isometry,
@@
$\mathbb{E}\left|\int_{0}^{T}\langle\Lambda^{2\beta}\varphi,P_{m(\varepsilon)}\,\mathrm{d}W(s)\rangle\right|^{2}\leqslant\|\Lambda^{2\beta}\varphi\|_{L^{2}({0,T};H)}.$
Hence we can pass to a further subsequence so that $\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi,P_{m(\varepsilon)}\,\mathrm{d}W(s)\rangle\rightarrow 0$, almost surely, for all $t\in[0,T]$. Finally, the convergence of the last term follows from the fact that $g\in L^{2}([0,T];H)$ and Hölder’s inequality. Therefore we conclude that, $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
@@
Step 2. $\bar{\theta}$ satisfies $H^{\alpha-2\beta}$-energy inequality. Applying Itô’s formula to $\bar{\theta}_{\varepsilon,m(\varepsilon)}$, we deduce that $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\quad\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (6.6)
$=\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle$
$+\frac{\varepsilon}{2}\left\|P_{m(\varepsilon)}\Lambda^{\alpha}\right\|_{HS}^{2}t+\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}g(s)\right\rangle\,\mathrm{d}s.$
Letting $\varepsilon\rightarrow 0$ in (6.6), it follows from the fact that $\bar{\theta}_{\varepsilon,m(\varepsilon)}\rightarrow\bar{\theta}\,\text{ in }\mathbb{X}_{\alpha,\beta}$ and the lower semi-continuity of the $H^{s}(\mathbb{T}^{2})$ norms that
$\quad\frac{1}{2}\|\bar{\theta}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\bar{\theta}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\leqslant\liminf_{\varepsilon\rightarrow 0}\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
Let $M^{\varepsilon,m(\varepsilon)}_{2}(t):=\int_{0}^{t}\sqrt{\varepsilon}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\,\mathrm{d}W(s)\right\rangle$. Applying Doob’s $L^{2}$-inequality, we obtain
$\mathbb{E}\left(\sup_{t\leqslant T}\left|M_{2}^{\varepsilon,m(\varepsilon)}(t)\right|^{2}\right)\leqslant 4\mathbb{E}\left\langle\langle M_{2}^{\varepsilon,m(\varepsilon)}\right\rangle\rangle_{T}\leqslant 4\varepsilon\mathbb{E}\int_{0}^{T}\left\|\bar{\theta}_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\rightarrow 0.$
Hence we can pass to a further subsequence so that $M_{2}^{\varepsilon,m(\varepsilon)}(t)\rightarrow 0$ almost surely. By the scaling regime (2.5), the correction term $\frac{\varepsilon}{2}\left\|P_{m(\varepsilon)}\Lambda^{\alpha}\right\|_{HS}^{2}t\rightarrow 0$. Finally, according to the weak convergence $\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m(\varepsilon)}\rightarrow\Lambda^{2\alpha-2\beta}\bar{\theta}$ in $L^{2}([0,T];H)$ and the strong convergence $P_{m(\varepsilon)}g\rightarrow g$ in $L^{2}([0,T];H)$, it holds that
$\left|\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m(\varepsilon)},P_{m(\varepsilon)}g\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta},g\right\rangle\,\mathrm{d}s\right|\rightarrow 0.$
@@
Step 3. $\bar{\theta}=\theta$ in $\mathbb{X}_{\alpha,\beta}.$ Thanks to the choice $\theta\in\mathcal{C}_{0}$, it follows from Definition 1.1 that $\theta=\bar{\theta}$ in $\mathbb{X}_{\alpha,\beta}$. Hence $\pi^{\varepsilon,m,\theta}$ converges to $\delta_{\theta}$ in law on $\mathbb{X}_{\alpha,\beta}$ as $\varepsilon\rightarrow 0$. ∎
@@
For every $\varepsilon,\delta(\varepsilon)>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta(\varepsilon)}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 1.2 with initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m(\varepsilon)}$ be the solution of (3.1) with initial data $\theta_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$. Suppose that $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1).
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta(\varepsilon)}$ be the law of $\theta_{\varepsilon,\delta(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (1.7) holds for $(\varepsilon,\delta(\varepsilon))$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
@@
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m(\varepsilon)}$ be the law of $\theta_{\varepsilon,m(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.5) holds for $(\varepsilon,m(\varepsilon))$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$, the restricted lower bound (6.7) holds as well.
@@
Proof of Proposition 6.1. For any open set $G\subset\mathbb{X}_{\alpha,\beta}$ and any $\eta>0$, there exists an element $\theta^{0}\in G\cap\mathcal{C}$ such that
@@
Using (6.8), (6.9), and the arbitrariness of $\eta$, we obtain that